From GlueXWiki
Jump to: navigation, search

November 21, 2013 FDC meeting


  1. Installation [1] (all)
  2. Engineering (Bill)
  3. Electronics (Chris, Nick)
  4. Mini-DAQ for FDC tests [3] (Beni)
  5. Other


Participants: Bill, Dave, Nick, Chris, Simon, Vlad, Beni, Elton, Hovanes (for the LV control discussions) and Lubomir.


- Dave: right now techs are working on the mesh, it will be finished by the end of the day or tomorrow. Tom prefers only after that the FDC to be moved on the platform (today or tomorrow morning). In the mean time Tom and Dave tested the cooling system for leakage with He at 10-15 psi using the big testing unit used for the magnet. A leak was found in the connection of one of the manifolds and Casey fixed it. Dave made the houses and the fittings needed to connect the FDC to the chiller once it is on the platform. As of yesterday we have also chilled water and the connections to the chiller are ready.

- We discussed details how to connect/bundle/route the LV cables. The cables will be laying on the floor all over the platform for some time and we want to make sure the cables and especially the connectors are at a safe place. Most likely we have to run the individual cables around the pair spectrometer magnet (these are 18m long cables) and then bundle them in groups of 8 or 9. The exact procedure will be clear once we start cabling. We should have then the bundles routed so that they don't interfere with the signal cables. We want to use regular cable ties at least close to the connectors (for strain relief), but then along the mesh we cab use Velcro ties (Nick has now narrower ties with bigger friction).


- The raw data from the last survey is now available (attached above, all the numbers are in meters). At a first look the thicknesses of the packages are OK, +/- 0.3mm from the nominal, but this requires more studies. The main question was if the prisms can be used to measure the z positions with a few hundred microns resolution, but the data we have are not enough to make a conclusion. Still one can see that the z positions of the prisms at one package vary by up to 6mm but Bill says this is possible given the way the prisms are glued to the gussets. The distances between the packages from the SMR and from the prism measurements are similar (up to few mm) for the package #1-#2 and #2-#3. There's about 4cm difference for the distance between #3 and #4, but there one has to take into account the SMR holder sizes. Certainly all this require careful analysis.

- Important is that the prisms were survey from a big distance of 6-8 m, chosen by the survey people to have very similar angles w.r.t. prism surfaces. However we will not have that space between the FDC in the magnet and the FCAL in the current position. Bill will talk to the survey group to find if they have estimated the effect of distance to the targets. After the meeting Bill talked to Jim Dahlberg: the prisms are accurate at 100 micron level if viewed at +/-5cm transversely at 3m from the target; at +/- 10cm the accuracy is 170 microns.

- Once the mesh is installed Bill wants to make sure CDC will fit. A wooden disk with bearings inserted in the rails will be used for that. The disk will be inside the mesh while cabling the FDC so that we can do checks from time to time. Bill thinks that at least initially the cables shouldn't be connected to the mesh very tightly so that we have the freedom to shape the mesh if it is not cylindrical.


- We have everything needed to start the cabling. According to Nick with the temporary 120V/20A power we can't have more than 4 LV modules in the crate, so we will be doing the LV tests in groups of several modules.


- Beni installed the computer "joebesser" on the platform close to the north racks. Hovanes set-up and tested the LV control system that can be used from "joebesser" under "halld" account. Since we have to change the LV module configuration and this requires ioc reboot, Hovanes gave us (at least temporary) privileges to do this.